# PersonaFlow: Boosting Research Ideation with LLM-Simulated Expert Personas

### v1

@@ 5.2.1. Patterns in Users’ Customization of LLM-Generated Personas
Figure 8. P20’s nodes flow. The user started by exploring different persona and RQ nodes broadly, and then decided to dive deep into interacting with one “remote” persona (“Spatial Gene Profiler”) with the topic “single-cell RNA sequencing and spatial transcriptomics.” Though the persona was not immediately related to the P20’s research, the recommended literature was surprisingly relevant to P20’s work and exactly what they read recently. This “synergy” quickly elicited P20’s interest in a deeper exploration of this persona.
5.2.2. Association between user behaviors during system use and overall perception of the ideation process.
In this section, we explore how the modifications and adaptations made by users to the personas influence the outcomes of the research questions. We analyzed the association between users’ behavior factors and their perception of the ideation process measured through ratings collected through the post-session survey. A set of linear regression analyses was conducted with post-session survey ratings as dependent variables and users’ behavioral factors (including the number of edits over RQ/critique/persona nodes and the percentage of nodes used by the user for subsequent generation) as predictors (independent variables). We measured the user’s percentages of nodes used by computing the percentage of nodes used for generating subsequent nodes out of all nodes generated. The regression analysis results are shown in Table 2.
More edits on the generated RQs correlated with their lower perceived reliance on the system. Our findings showed that users’ number of edits on RQ nodes was negatively associated with the overall perceived creativity of RQ nodes ($t(36)=-2.060,p=0.039^{*}$), the perceived extent to which the system changed their way of thinking ($t(36)=-3.548,p<0.001^{***}$) and reliance on the system ($t(36)=-3.164,p=0.002^{**}$). The number of users’ edits over RQ nodes was negatively associated with the creativity of RQs, their perceived extent to the system changed their ways of thinking, as they identified the need to modify RQ nodes when they felt that the generated research questions were inadequate or misaligned with their ideation goals. Thus by allowing users to make edits over the RQ nodes, the system provided users with greater control over the content, which consequently led to a lower perceived reliance on the system.
More exploration improved user-perceived quality of generated critique and research outline. We found significantly positive associations between users’ percentages of nodes used (of Nodes Used) and their ratings given towards the helpfulness of generated critiques (Critique Helpfulness) during the post-session survey ($t(36)=3.070,p=0.002^{**}$). A similar positive association also exists between the total number of nodes created by each user (Total Node Count) and their critique helpfulness (Critique Helpfulness) survey ratings ($t(36)=3.639,p<0.001^{***}$). The results indicate that as users explored more nodes, they received higher-quality critiques from personas. As mentioned by the participants, new insights were discovered through reading critiques during the exploration. P13 mentioned that, “the system provided a new approach for exploration that I hadn’t initially considered,” emphasizing the importance of diverse exploration in enhancing the quality of user-perceived critiques.
More edits on personas improve users’ perceived recall of the generated ideas and perceived control during use. More edits made on personas (# of edits on Persona nodes) led to users’ enhanced recall (Recall of Ideas) of the generated ideas ($t(36)=2.830,p=0.005^{**}$), As shown in Table 2. Users’ perceived control of the system (Perceived Control) was also improved through more edits over personas (# of edits on Persona nodes) during the exploration ($t(36)=2.617,p=0.009^{**}$). The perceived relevance and helpfulness of critiques did not show a significant correlation with users’ persona edits. While we cannot state that persona edits enhance ideation outcomes, our findings suggest that utilizing a higher percentage of nodes contributes to fostering a stronger sense of recall over the ideas co-created between the user and the system and allows users to have a stronger sense of agency in shaping the generated ideas. Some participants also mentioned the mind-map-based design specifically helped them identify and recall system-generated research ideas, “… I personally also create workflows to have a clear direction (during research) or just brainstorming … it helped me recall and identify the ideas I had during the process … ” as noted by P9.
Figure 9. P9’s node flow (displaying 22 out of 34 nodes). The participant employed a divergent exploration strategy when creating the flow. The participant explored literature generated by different personas and appreciated the perspectives from different research communities (i.e., privacy, accessibility and AI Ethics research).
5.3. Case studies: Exploring Diverse Use Cases with Multi-Personas (RQ3)
After the task-based study session, we invited participants to an optional 30-minute free exploration session where they freely used the system’s features without guidance or constraints. Researchers observed their usage of different nodes. This section analyzes their behavior and think-aloud data (N=10) and presents cases demonstrating how they collaborated with AI to co-create research ideas and questions, reflecting a mixture of divergent or convergent thought processes.
5.3.1. From General to Specific: How Persona Literature Recommendations Shaped a Targeted Research Path
Figure 8 shows an example of participant P20 generated a flow with a depth of six iterations. The participant first examined all personas generated using P20 initial research question. The user chose to explore personas closely aligned with P20’s own expertise. For example, a persona “Bio Data Interpreter” quickly caught the user’s attention, as the persona recommended relevant literature to analyzing biological data for cancer research. P20 expressed that (“Some of the papers are what I just read, exactly the same paper.. that’s pretty interesting”). During the first three rounds of exploration, the user chose one of the personas “Spatial Gene Profiler.” The persona was beyond the user’s immediate research interest, but after reviewing the literature suggested by the persona, the user found expected connections between the persona’s suggestion and P20 initial interest (“Some of the papers I didn’t expect to be relevant actually linked back to my research. I was pretty surprised that it connected things I hadn’t thought of before…”). The user’s choice of this persona, driven by the interest in P20’s proposed literature, marked a shift in P20’s direction of exploration from broadly examining different personas to diving deep into a specific research direction. The topic of exploration, originally the broader topic of “spatial distribution of RNA in cells,” shifted to focus on ”single-cell RNA sequencing and spatial transcriptomics.” The user then subsequently generated three more iterations of research questions branching from this chosen persona of “Spatial Gene Profiler.” Looking at the long exploration journey, the user suggested us to include a feature for a quicker navigation back to the previous nodes, especially for easy revisiting of the key RQ nodes or literature summaries: “If I want to review some previous results, it takes time to scroll back… it takes time to like for me to drop back and revisit that.”
Figure 10. P15’s node flow (displaying 9 out of 46 nodes). The participant utilized literature nodes to generate different nodes, e.g., persona nodes and critique nodes, during the free exploration session.
5.3.2. Research Question Refinement through Persona Feedback: Confidence and New Directions
Figure 9 presents a case of P9 exploring all critiques brought by the persona “Privacy and Security Expert” and requesting to generate RQ nodes for each of the critiques before deciding which one of them to proceed with. The participant started by exploring the initial research question of “How do blind users use Gen AI tool to manage privacy in images and videos and for what purpose?” When browsing different system-generatRQs, P9 mentioned seeing multiple RQs similar to their own idea in mind helped improve their own confidence in further exploring the topic, “I really thought about similar research questions to explore, but I was not sure, should I go for it or not… what really helps is that the system confirmed my thoughts, seeing the RQs that are similar to my own idea… ,” as P9 mentioned when reflecting on the ideation outcomes generated by the system. The participant eventually chose the RQ “What are the experiences and challenges faced by blind users when co-designing Gen AI tools for managing privacy in images and videos?”, because it provided new insights that they had not considered before. P9 mentioned, “This set of research questions from technical feasibility… then privacy, risk, and security… but user design… this co-designing ideas, those I haven’t really thought before. In that sense, it’s most appealing to me.” During the exploration, the participant also shared their thoughts on appreciating the critiques and suggested literature from different personas who were not in P9’s research community. More specifically, the literature topic of “Challenges Blind Users Face In AI Tool Design” proposed by the persona “Accessible Technology Researcher” introduced a new literature space in addition to the ones introduced by a previous persona “Privacy and Security Expert.”
5.3.3. Dynamic Node Connections: Exploring Literature and Critiques through Persona-Driven Pathways
Figure 10 presents a case where P15 varied the connections between different nodes to explore different thought processes. For example, users can request PersonaFlow to generate multiple persona nodes from a specific recommended literature or request a persona to continue generating critiques from a set of literature. In the example case, P15 applied both pathways during the free exploration session. P15 thought aloud that the functionality of generating literature-sourced personas was helpful in identifying additional literature from a broad set of topics, while generating critiques earlier during the exploration helped confine the literature space. Specifically, in Figure 10, P15 chose the persona “AML Stem Cell Specialist”, as noted “… unlike Google Scholar, you need a clear idea from the start to find specific papers… the generated persona (of PersonaFlow) probes my interest… .”
5.3.4. Adapting Research Questions through Persona Feedback: A Case Study of RQ Re-framing.
Figure 11 depicted an example where the participant went back and re-framed their original RQs after evaluating the personas and critiques, to better align the RQs more closely with their evolving exploration interests. This was a common observation among the participants. Participant P5 first generated three personas from their initial RQ node. After reviewing the literature nodes, the user found them-self interested in none of the suggested literature. As a result, the participant revised the original research question from “How to develop a wearable assistant on AR smart glasses to help UX research conduct & record field study or shadowing study?” to “How to develop a wearable AI LLM (Large language model) assistant on AR smart glasses to help UX research conduct & record field study or shadowing (experimenter observe the participants in shadow) study?” The revision led to three additional personas, two of which the participant found helpful, prompting further exploration through follow-up literature suggestions. Specifically, the participant found the newly generated persona “UI Research Expert,” with experience in “field studies and user shadowing,” particularly helpful in providing relevant literature. Additionally, the persona “AI LLM Expert,” specializaing in “designing user-friendly interfaces for AR devices,” closely aligned with their exploration interests.
Figure 11. P5’s node flow (displaying 15 out of 63 nodes). The participant revised their initial research question after reviewing the generated personas and literature retrieved based on the personas.
@@ 6. Discussion
6.1. Empowering Interdisciplinary Research Ideation through Customizable Multi-Persona AI: Enhancing User Agency, Idea Ownership, and Knowledge Acquisition
@@
Evaluating the PersonaFlow system with researchers showed that its multi-persona design enhances research ideation by improving outcomes without adding cognitive load (section 5.1). Beyond past studies focusing on the perspective of audience (Petridis et al., 2023), we demonstrated that incorporating expert AI personas with domain-specific expertise fosters interdisciplinary research. Users effectively expressed different thinking patterns and addressed diverse use cases (section 5.3), with the help of the interactive mind map and easy persona customization, especially when refining ideas through iterative revisions. The system also benefits users exploring less familiar domains by providing domain-specific concepts and theories (section 5.2.2). For researchers new to a field, challenges include knowing the right questions to ask or bridging communication gaps (August et al., 2024; MacLeod, 2018, 2018). In the context of the PersonaFlow system, the interaction between a user and AI-simulated expert personas offers instant, iterative feedback and scaffolding to key concepts, which in turn, improves knowledge acquisition and facilitates future communication with human experts.
Additionally, our findings (section 5.1) indicated the design of supporting users’ edits over AI-generated expert personas enhanced their perceived control over the system, likely due to reduced cognitive load and increased user autonomy By further examining users’ behaviors during persona customization (section 5.2), we identified that the uniqueness of a persona profile as an interface of communication with the system has provided users with a means to project their own intentional states (Dennett, 1971, 1987) onto the AI, often with the intention of obtaining corresponding informational responses from the AI system. This design promotes user agency while maintaining ideation support drawing from the knowledge from different domains. This can also potentially improve users’ trust in the persona and the system’s subsequently generated outcomes while reducing their cognitive load (De Graaf and Malle, 2017). Prior work (Guo et al., 2024) has raised concerns of reduced idea ownership during brainstorming using AI for creativity support, while our work demonstrates that allowing users to modify AI-generated personas can potentially alleviate this issue as a design that also improves the transparency and explainability of the system.
@@
6.2. Navigating Cognitive Biases in Human-AI Persona Interaction: Confirmation Bias and Social Conformity
Cognitive biases have been a prevalent concern in human-AI co-creation systems (Bertrand et al., 2022). Although the findings demonstrated the utility of the system in providing research ideation support, we noted potential biases when users interacted with the personas during system use. In this section, we discuss what the biases might be and how they might lead to a sub-optimal ideation process.
Confirmation biases in selecting and editing personas. Participants exhibited potential confirmation biases (Rastogi et al., 2022) when generating follow-up nodes during the system use, which reflects biases from both human and AI sources in co-creation systems (Buschek et al., 2021). More specifically in PersonaFlow, human bias arises from users’ decisions in selecting and editing personas, while AI bias reflects inherent biases in the LLM’s training data.
Recent research indicates that users’ backgrounds and preferences contribute to confirmation biases in human-AI co-creation (Liu et al., 2024). Users’ affinity for AI personas close to their background was also suggested through our findings (section 5.2.1), including the tendency to edit personas by injecting their own background or expertise. Users’ trust increased when they recognized research questions similar to their ideas or when the system recommended their own works. While this enhances trust, it may reinforce confirmation biases, limiting the diversity and novelty of insights. This aligns with Cognitive Dissonance Theory (Braga, 1972; Harmon-Jones and Mills, 2019), suggesting individuals prefer assistance aligning with their beliefs. Our observations further echo this phenomenon.
We speculate that guiding users to “choose” from generated expert personas exacerbates these biases, as it is less cognitively demanding than creating or critically editing personas. To mitigate this, future designs should encourage critical evaluation and integration of diverse persona traits possibly by reducing cognitive load and offering additional support. For instance, the system can provide suggestions of potential traits from domains beyond the user’s expertise when the user is editing or selecting personas.
Social conformity to personas with authority leading to collaborative fixation. Another potential cognitive bias we noticed was social bias, specifically authority bias (Milgram, 1963; Juárez Ramos, 2018) and social conformity (Stallen and Sanfey, 2015). Our findings (section 5.2.1) showed that users prefer interacting with personas of stronger perceived expertise, even manually editing them to enhance their authority. When users edit persona profiles, they may not have introduced new information, but aim to enhance their trust in the persona. This behavior reflects informational conformity (Stallen and Sanfey, 2015), where individuals align with perceived authority figures, assuming they possess greater knowledge.
This can be problematic because it can result in an over-reliance on the system due to a perceived, but misleading, sense of authority. As a consequence, it may overshadow valuable contributions from less authoritative but equally informative personas or sources. Combined with the anchoring effect (de Wilde et al., 2018) during the initial choice of persona, these behaviors could lead to collaborative fixation (Kohn and Smith, 2011), constraining the exploration of alternative ideas. Our speculation implies that the design of persona-based human-AI collaboration systems needs to account for social cognitive biases to ensure diversity in terms of the information users use to validate and establish accurate alignment of expectations on AI personas’ reliability of the provided information.
@@ 6.3. Design Implications
6.3.1. Facilitating diverse research ideation and scaffolding through AI-generated personas.
Leveraging the fast and adaptive responses of LLMs, prior studies have explored their effectiveness for idea iteration, scaffolding, and immediate feedback (Dhillon et al., 2024; Heyman et al., 2024). During free exploration sessions (section 5.3), users’ behaviors of using multiple personas to explore the literature spaces from different research communities suggest the design of research-oriented personas should emphasize the common values and norms of different communities. Future designs should conduct domain-specific literature surveys to synthesize personas that capture unique methodologies and theoretical perspectives. Participants (P15, P3) also highlighted the importance of selecting different types of literature resources at each stage of the research ideation process, considering the background of the researcher. More specifically, for earlier-stage researchers, the system should provide an overview and survey, while for more advanced users detailed methodologies, theories, and latest developments in the field would be more beneficial. Potential designs for personalizing the ideation process, such as literature recommendations based on semantic similarity between retrieved papers and users’ own publications, can be employed to enhance the relevance with awareness of users’ current stages of research. Features such as incorporating predicted user intent and background as contextual information (Pfautz et al., 2015), to support the establishment of the mutual theory of mind (MToM) (Wang et al., 2021) would increase the utility of the system catering to different user groups.
@@ 7. Limitation and Future Research
While our study demonstrates the potential of using LLM-simulated expert personas to facilitate the research ideation process, several limitations should be noted. We did not conduct systematic evaluation of the quality and accuracy of the personas that the system generated nor the literature they recommended. The personas’ knowledge might inherently be limited by the dataset on which large language models were trained on when generating search queries. Another limitation of our study is the absence of a direct comparative baseline with similar ideation tools. While there are currently no ideation instruments that closely match the functionality of PersonaFlow, this limits our ability to objectively compare its usefulness. Future research should consider evaluating possible benchmarks or developing simplistic frameworks for comparative analysis. Third, while PersonaFlow supports visual programming-styled interactions, which is recognized for its potential to improve user engagement and usability (Resnick et al., 2009; Wu et al., 2022), our study did not explicitly assess the effectiveness of this functionality through the design of controlled study. Future studies should examine the influence of these visual interactions on user experience and ideation results. Last but not least, our study does not address two important aspects: system interactions outside direct persona manipulation and social interactions other than solicitation of feedback and literature information, such as natural conversations. Future research should explore the extent to which these social elements genuinely impact the ideation process versus merely reflecting users’ inherent biases or psychological tendencies, potentially through controlled experimental designs that isolate the effects of social traits against informational traits.

### v2

@@ 6. Discussion
6.1. Bridging Interdisciplinary Research Ideation and User Agency through Customizable Multi-Persona AI
@@
Evaluating the PersonaFlow system with researchers showed that its multi-persona design enhances research ideation without adding cognitive load (section 5.1) over the use of singular personas. Beyond past studies focusing on the perspective of audience (Petridis et al., 2023), we demonstrated that incorporating expert AI personas with domain-specific expertise fosters interdisciplinary research. Users effectively expressed different thinking patterns and addressed diverse use cases (section 5.2), with the help of the interactive mind map and easy expert profile customization, especially when refining ideas through iterative revisions. The system also benefits users exploring less familiar domains by providing domain-specific concepts and theories (section 5.3.1), providing instant feedback and scaffolding for knowledge acquisition. Additionally, allowing users to edit AI-generated personas enhanced their perceived control and autonomy (section 5.1). For researchers new to a field, challenges include knowing the right questions to ask or bridging communication gaps (August et al., 2024; MacLeod, 2018, 2018). In the context of the PersonaFlow system, the interaction between a user and AI-simulated expert personas offers instant, iterative feedback and scaffolding to key concepts, which in turn, improves knowledge acquisition and facilitates future communication with human experts. our findings (section 5.1) indicated the design of supporting users’ edits over AI-generated expert personas enhanced their perceived control over the system, likely due to reduced cognitive load and increased user autonomy. By further examining users’ behaviors during expert profile customization (section 5.2), we identified that the uniqueness of an expert profile as an interface of communication with the system has provided users with a means to project their own intentional states (Dennett, 1971, 1987) onto the AI, often with the intention of obtaining corresponding informational responses from the AI system. This design promotes user agency while maintaining ideation support, drawing from the knowledge from different domains, meanwhile improving trust, reducing cognitive load (De Graaf and Malle, 2017), and addressing concerns about reduced idea ownership (Guo et al., 2024), as modifying expert profiles helps users retain ownership and better understand the system’s capabilities. This can also potentially improve users’ trust in the persona and the system’s subsequently generated outcomes while reducing their cognitive load (De Graaf and Malle, 2017). Prior work (Guo et al., 2024) has raised concerns of reduced idea ownership during brainstorming using AI for creativity support, while our work demonstrates that allowing users to modify AI-generated personas can potentially alleviate this issue as a design that also improves the transparency and explainability of the system.
6.2. Cultivating Critical Thinking Through Multi-Perspective Expert Simulation
Our findings indicate that the multi-perspective design of PersonaFlow not only supports interdisciplinary exploration but also triggers higher-order thinking processes crucial for research ideation. In particular, we observed users repeatedly engaging in core critical thinking elements defined by past research (Facione, 1990). PersonaFlow’s design of simulating experts tends to incur participants’ activities related to self-regulation and evaluation by driving users to think about the alignment of background and interest between themselves and system-generated expert profiles. We observed participants re-stating the scope of their own research domains and clarifying terminology when browsing various expert profiles. For instance, by scanning expert profiles, some users reformulated their original research questions in more precise terms (e.g., adding the keyword “LLM” to reflect their actual research goal). Some instances of evaluation (e.g., where the user identifies an expert profile to have a similar background to theirs) were also found to be associated with confirmation of users’ pre-existing research ideas and consolidation of their trust and willingness to further engage with the expert perspective, contributing to trust-building but also raising the risk of over-reliance on AI. Additionally, the system’s mind-map structure and iterative interactions promoted user-led reflection, through activities of inference from newly suggested ideas or concepts by different experts during the exploration. Participants were observed to frequently speculate on new connections or follow-up topics based on expert suggestions (e.g., pivoting from a broad question about “spatial distribution of RNA” to a narrower focus on single-cell transcriptomic). Our findings suggest that the multi-perspective interaction design appears to contribute to users’ interdisciplinary research ideation process, by promoting different levels of critical thinking activities introducing both new and more specific research topics and directions. Such observations highlight the potential application of using AI-simulated experts for fostering higher-order critical thinking skills in interdisciplinary research and education (Halpern, 1998; Lee et al., 2024). On the other hand, we observed limited instances of users’ explanation activities beyond what was prompted by the think-aloud protocol. This is potentially because PersonaFlow did not design features to support or stimulate explicit rationalization of their behavior and decisions. In future iterations, designers could explore using AI-simulated personas that pose reflective questions (Kumar et al., 2024), prompting users to elaborate further on their thought processes and rationales.
@@ 6.2. Cultivating Critical Thinking Through Multi-Perspective Expert Simulation
6.3. Navigating Cognitive Biases in Human-AI Persona Interaction: Confirmation and Social Biases
Cognitive biases have been a common concern in human-AI co-creation systems (Bertrand et al., 2022). Although the findings demonstrated the utility of the system in providing research ideation support, we noted potential biases when users interacted with the expert personas during system use. In PersonaFlow, confirmation bias arose from both users’ expert choices and the LLM’s inherent biases (Rastogi et al., 2022; Buschek et al., 2021). Participants tended to select or edit personas with backgrounds aligning with their own (section 5.3.5), reinforcing existing beliefs (Liu et al., 2024). They also trusted personas more when they aligned with personal experiences, potentially limiting ideation diversity. This aligns with Cognitive Dissonance Theory (Braga, 1972; Harmon-Jones and Mills, 2019), suggesting individuals prefer assistance aligning with their beliefs. Our observations further echo this phenomenon. Another potential cognitive bias we noticed was social bias, specifically authority bias (Milgram, 1963; Juárez Ramos, 2018) and social conformity (Stallen and Sanfey, 2015). Users favored expert personas perceived as more authoritative (section 5.3.5), even modifying profiles to enhance perceived expertise. This can lead to over-reliance on the system and overshadow less authoritative but equally informative expert personas. Combined with the anchoring effect (de Wilde et al., 2018), these factors risk collaborative fixation (Kohn and Smith, 2011), constraining alternative perspectives.
@@ 6.4. Design Implications
6.4.2. Beyond Research Ideation: Facilitating Diverse Scaffolding Through AI-Simulated Personas
Leveraging the fast and adaptive responses of LLMs, prior studies have explored their effectiveness for idea iteration, scaffolding, and immediate feedback (Dhillon et al., 2024; Heyman et al., 2024). While our study focused on supporting the research ideation process, the core design principle of using customizable AI personas can be extended to a variety of other fields such as design, consulting, content creation, product development, and decision-making in strategic planning. During free exploration sessions (section 5.2), participants used multiple personas to examine literature from various research communities. This behavior suggests that personas, whether intended for research or other domains, should highlight the shared values, norms, and conceptual frameworks typical of those communities. For instance, service designers (Han et al., 2018) or consulting teams could develop customizable personas representing different user groups to surface their diverse perspectives and values.
Additionally, participants (P15, P3) emphasized the importance of adapting literature resources (e.g., foundational concepts vs. cutting-edge methods) to a researcher’s prior knowledge. Incorporating predicted user intent and background as contextual information (Pfautz et al., 2015) and supporting the establishment of a mutual theory of mind (MToM) (Wang et al., 2021) can help the system deliver content at the appropriate level of detail. Analogously in other domains, newcomers may require broader overviews, while experts benefit from deep-dives into specialized methodologies or emerging trends. Future systems could personalize system outputs through the interface of simulated personas, such as adopting the complexity of outputs, based on user’s current knowledge or project scope conveyed through customizable persona profiles.
Our findings also suggest that the persona customization interface offers users a stronger sense of ownership towards co-creation outcomes, helping them to more intuitively shape AI-driven feedback toward their immediate needs. Beyond research, this same principle can foster rapid ideation in general. For instance, an online content creation team might instantiate multiple personas representing different audience segments, and iterate on their profiles to better align with shifting project requirements or newly discovered audience insights. By allowing fine-grained control over persona attributes, systems can encourage diverse and adaptable exploration of multiple perspectives and critical thinking.
@@ 7. Limitation and Future Research
While our study demonstrates the potential of using LLM-simulated expert personas to facilitate the research ideation process, several limitations should be noted. We did not conduct systematic evaluation of the quality and accuracy of the personas that the system generated nor the literature they recommended. The personas’ knowledge might inherently be limited by the dataset on which large language models were trained on when generating search queries. Another limitation of our study is the absence of a direct comparative baseline with similar ideation tools. While there are currently no ideation instruments that closely match the functionality of PersonaFlow, this limits our ability to objectively compare its usefulness. Future research should consider evaluating possible benchmarks or developing simplistic frameworks for comparative analysis. Last but not least, our study does not address two important aspects: system interactions outside direct persona manipulation and social interactions other than solicitation of feedback and literature information, such as natural conversations. Future research should explore the extent to which these social elements genuinely impact the ideation process versus merely reflecting users’ inherent biases or psychological tendencies, potentially through controlled experimental designs that isolate the effects of social traits against informational traits.